Consider the experiment of selecting a card from an ordinary deck of 52 playing cards and determine the probability of the stated event.

A card with an odd number and a prime on it is drawn.

Found 2 solutions by edjones, drk:
Answer by edjones(8007)

Probability a 3, 5 or 7 is drawn is 12/52 = 3/13

Answer by drk(1908)

Lets assume that ace = 1 and jack, queen, and king have no numbers and cannot be expressed using numbers.
The odd primes are: hearts - 2,3,5,7; diamonds - 2,3,5,7; club - 2,3,5,7;
spade - 2,3,5,7
P(odd and prime): 16/52 = 8/26 = 4/13